The annual sales for Jones, Inc. were $4.62 million last year. The firm's end-of-year balance sheet was as follows: Current assets: $ 505,000 Liabilities: $992,000 Net fixed assets: $1,479,000 Owners equity: $992,000 Total assets: $1,984,000 Total: $1,984,000 Jones income statement for the year was as follows: Sales $4,620,000 Less: cost of goods sold (3,492,000) Gross profit $1,128,000 Less: operating expenses (501,000) Net operating income $627,000 Less: interest expense (92,000) Earnings before taxes $535,000 Less: taxes (35%) (187,250) Net income $347,750 a. Calculate Jones total asset turnover, operating profit margin, and operating return on assets. b. Jones plans to renovate one of its plants and the renovation will require an added investment in plant and equipment of $1.09 million. The firm will maintain its present debt ratio of 50% when financing the new investment and expects sales to remain constant. The operating profit margin will rise to 13.8%. What will be the new operating return on assets ratio for Jones after the plant's renovation? c. Given that the plant renovation in part (b) occurs and Jones interest expense rises by $53,000 per year, what will be the return earned on the common stockholders' investment? Compare this rate of return with that earned before the renovation. Based on this comparison, did the renovation have a favorable effect on the profitability of the firm?
Answer #1

(A)
Asset turnover = Sales/ Assets = $4620000/(1479000 + 505000) = 2.33
operating profit margin = Operating Income / Sales = 627000 / 4620000 = 13.57%
operating return on assets = Operating Income / Assets = 627000 / (1479000 + 505000) = 31.60%

(B)
new assets = Current Total Assets + New Assets = 1984000 + 1090000 = 3074000
Sales = 4620000 (will remain same)
Operating margin = 13.80%
Operating income = Sales * 13.80% = 4620000 * 13.80% = $637560
New operating return on assets = New Operating Income / Assets = 637560 / 3074000 = 20.74%

(C)
Current net income = $347750
New Net income = [New Operating Income - new interest] * (1-tax rate)
= [637560 - (92000 + 53000)]*(1-35%)
New Net Income= $320164

Current equity = $992000
New equity = $992000 + 1090000*0.50 = $1537000
{half of new loan financed through debt and rest through equity}

old ROE = 347750/992000 = 35.06%
new ROE = 320164/1537000 = 20.83%

The Rate of return has been decreased due to increase in equity.